When A = 1, B = 0, C = 1. Then:

F = C + C'B + BA' = 1


1
Expert's answer
2021-04-15T07:21:15-0400

"A = 1; B = 0;C = 1"

"C' = 0;A'=0"

"C'B = 0*0=0;BA'=0*0=0"

"F = C + C'B + BA'=1+0+0=1"

"\text{the statement is correct}"